OpenStudy (anonymous):

Help! How to make an extraneous equation? I need to make an equation that when substituted into it's original equation, doesn't equal the original answer (I hope that makes sense) and the equation I made is non extraneous. My equation was the square root of (x+9) plus 2 = 20. But, it was non extraneous.

OpenStudy (anonymous):

try \[\sqrt{x+9}+2=1\]

OpenStudy (anonymous):

That worked, thanks! to be sure, is the solution x=-7?

OpenStudy (anonymous):

and when I tried to see if it was extraneous, I got 4=-1

OpenStudy (anonymous):

there is no solution, but i guess the "extraneous" one would be \(-8\)

OpenStudy (anonymous):

Sorry, I'm confused. I solved the equation and got -7 haha

OpenStudy (anonymous):

\[\sqrt{x+9}+2=1\\ \sqrt{x+9}=-1\\ x+9=1\\ x=-8\]

OpenStudy (anonymous):

oh ok! I put -7 instead of 8 for some reason. Thanks!

OpenStudy (anonymous):

yw